# Motion Planning Diffusion: Learning and Adapting Robot Motion Planning with Diffusion Models

### v1

@@ V Limitations and Future Work
To incorporate first-order gradients while sampling from the posterior distribution, we use the approach from classifier-guided diffusion. Instead, we can use a projection step at each denoising step, such that the sample lies on some manifold (for instance, ensuring safety constraints) [112]. However, the projection can be computationally expensive, and its study is left for future work.
In this work, we left the trajectory duration unchanged since optimizing it during the denoising process is not trivial, as the phase-time derivative would constantly change. Thus, the joint velocity and acceleration limits are not constant in phase space. A research direction would be to optimize the trajectory duration, possibly using a separate B-spline curve to represent the phase-time derivative from eq. 23 [93].
The current inference speed highly depends on the current Python implementation. One common appointed disadvantage of diffusion models is that they are slow to sample due to the number of denoising steps. However, in our problem, much computation time is used for the cost function gradients. Even though single-cost computation is not slow (done at $\approx 100$Hz) and parallelizable, the costs must be computed sequentially due to the Python GIL. Using other programming frameworks, such as JAX [113], which uses JIT compilation, might be an alternative to improve speed further, which could lead to applying MPD in a model predictive control setting.
@@ Appendix A Hyperparameters Training
Table II shows the hyperparameters used for training the diffusion model in each task. We generate one trajectory per context (start and goal configuration), so the dataset size equals the total number of contexts and trajectories. The number of control points is task-dependent and chosen such that after fitting a $5$th order B-spline, we obtain over $99\%$ of collision-free paths with a minimal collision rate. The U-Net input dim is the number of dimensions the control points are projected at the network input. The number of entries in the U-Net dim multiplier is the U-Net depth (layers) ($3$ or $4$, depending on the task), while the numbers represent the channels in each layer of the U-Net, which are multiplied by the input dimension. Context out dim is context network output size. We use a fixed number of $N=100$ diffusion steps for all tasks. The context and the denoising networks’ parameters are optimized using mini-batch gradient descent with the Adam optimizer [108] and a learning rate ${3\times 10^{-4}}$. The $2$D models took approximately $12$ hours to train, and the more complex ones $24$ hours (these can be improved by early stopping). Large batch sizes are beneficial when training diffusion because if the dataset has $D$ points, the model is trained for $O$ optimization steps with a batch size $B$, then each data point should be seen roughly $OB/D$ times (approximately the number of epochs). The diffusion loss function eq. 7 includes an expectation over the $N$ diffusion time steps. Therefore, the pair of data (control points, diffusion step) is roughly seen $OB/(DN)$ times, which motivates using a larger batch size and a larger number of optimization steps than other generative models. The CVAE baseline model is trained with the same hyperparameters, as we did not find significant differences in performance.
@@ Appendix B Hyperparameters Inference
[Uncaptioned image] João Carvalho is a Ph.D. candidate at the Intelligent Autonomous Systems Lab of the Technical University of Darmstadt. He previously obtained his M.Sc. degree in Computer Science from the University of Freiburg (Germany). His research interests are centered around learning approaches for robot manipulation, from generative models for robot motion planning, imitation learning, and grasping.
@@
[Uncaptioned image] Piotr Kicki Piotr Kicki is a post-doc in the robotics team at IDEAS NCBR and an assistant professor at the Institute of Robotics and Machine Intelligence at Poznan University of Technology. He received his B.Eng. and M.Sc. degrees in automatic control and robotics from Poznan University of Technology, Poland in 2018 and 2019, respectively. He completed his Ph.D. from the same university in 2024. His main research interests include motion planning and applications of machine learning in robotics.
[Uncaptioned image] Dorothea Koert Dorothea Koert is an Independent Research Group Leader of the interdisciplinary BMBF junior research group IKIDA, which started in October 2020. She completed her Ph.D. from the Technical University of Darmstadt in February 2020. In 2019 she was awarded the AI-Newcomer award by the German Society for Computer Science (GI). During her Ph.D. she has worked on imitation learning and interactive reinforcement learning, for autonomous and semi-autonomous acquisition of motion skill libraries in human-robot collaboration.
[Uncaptioned image] Jan Peters Jan Peters is a full professor (W3) for Intelligent Autonomous Systems at the Computer Science Department of the Technical University of Darmstadt, deptartment head of the research department on Systems AI for Robot Learning (SAIROL) at the German Research Center for Artificial Intelligence (Deutsches Forschungszentrum fur Künstliche Intelligenz, DFKI), and a founding research faculty member of The Hessian Center for Artificial Intelligence. He has received the Dick Volz Best 2007 US Ph.D. Thesis Runner-Up Award, RSS - Early Career Spotlight, INNS Young Investigator Award, and IEEE Robotics & Automation Society’s Early Career Award, as well as numerous best paper awards. He received an ERC Starting Grant and was appointed an IEEE fellow, AIAA fellow, and ELLIS fellow.

### v3

@@ V Limitations and Future Work
To incorporate first-order gradients while sampling from the posterior distribution, we use the approach from classifier-guided diffusion. Instead, we can use a projection step at each denoising step, such that the sample lies on some manifold (for instance, ensuring safety constraints) [112]. However, the projection can be computationally expensive.
In this work, we kept the trajectory duration fixed, since optimizing it during the denoising process is not trivial, as the phase-time derivative would constantly change. Thus, the joint velocity and acceleration limits are not constant in phase space. A research direction would be to optimize the trajectory duration, possibly using a separate B-spline curve to represent the phase-time derivative from eq. 23 [93].
The current inference speed highly depends on the current Python implementation. One common appointed disadvantage of diffusion models is that they are slow to sample due to the number of denoising steps. However, in our problem, much computation time is used for the cost function gradients. Even though single-cost computation is not slow (done at $\approx 100$Hz) and parallelizable, the costs must be computed sequentially due to the Python GIL. Using other programming frameworks, such as JAX [113], which uses JIT compilation, might be an alternative to improve speed further.
@@ Appendix A Hyperparameters Training
Table II shows the hyperparameters used for training the diffusion model in each task. We generate one trajectory per context (start and goal configuration), so the dataset size equals the total number of contexts and trajectories. The number of control points is task-dependent and chosen such that after fitting a $5$th order B-spline, we obtain over $99\%$ of collision-free paths with a minimal collision rate. The U-Net input dim is the number of dimensions the control points are projected at the network input. The number of entries in the U-Net dim multiplier is the U-Net depth (layers) ($3$ or $4$, depending on the task), while the numbers represent the channels in each layer of the U-Net, which are multiplied by the input dimension. Context out dim is context network output size. We use a fixed number of $N=100$ diffusion steps for all tasks. The context and the denoising networks’ parameters are optimized using mini-batch gradient descent with the Adam optimizer [108] and a learning rate $3\times 10^{-4}$}. The 22D models took approximately $12$ hours to train, and the more complex ones $24$ hours (these can be improved by early stopping). Large batch sizes are beneficial when training diffusion because if the dataset has $D$ points, the model is trained for $O$ optimization steps with a batch size $B$, then each data point should be seen roughly $OB/D$ times (approximately the number of epochs). The diffusion loss function eq. 7 includes an expectation over the $N$ diffusion time steps. Therefore, the pair of data (control points, diffusion step) is roughly seen $OB/(DN)$ times, which motivates using a larger batch size and a larger number of optimization steps than other generative models. The CVAE baseline model is trained with the same hyperparameters.
@@ Appendix B Hyperparameters Inference
[Uncaptioned image] João Carvalho is a Postdoctoral Researcher at the Intelligent Autonomous Systems Lab of the Technical University of Darmstadt, where he obtained his Ph.D. in January 2025. He previously obtained his M.Sc. degree in Computer Science from the University of Freiburg. His research interests are centered around learning approaches for robot manipulation, from generative models for robot motion planning, imitation learning, and grasping.
@@
[Uncaptioned image] Piotr Kicki is a post-doc in the robotics team at IDEAS and an assistant professor at the Institute of Robotics and Machine Intelligence at Poznan University of Technology. He received his B.Eng. and M.Sc. degrees in automatic control and robotics from Poznan University of Technology, Poland in 2018 and 2019, respectively. He completed his Ph.D. from the same university in 2024. His primary research interests center on the application of machine learning to improve robot motion planning and control.
[Uncaptioned image] Dorothea Koert is an Independent Research Group Leader of the interdisciplinary BMBF junior research group IKIDA, which started in October 2020. She completed her Ph.D. from the Technical University of Darmstadt in February 2020. In 2019 she was awarded the AI-Newcomer award by the German Society for Computer Science (GI). During her Ph.D. she has worked on imitation learning and interactive reinforcement learning, for autonomous and semi-autonomous acquisition of motion skill libraries in human-robot collaboration.
[Uncaptioned image] Jan Peters is a full professor (W3) for Intelligent Autonomous Systems at the Computer Science Department of the Technical University of Darmstadt, department head of the research department on Systems AI for Robot Learning (SAIROL) at the German Research Center for Artificial Intelligence, and a founding research faculty member of The Hessian Center for Artificial Intelligence. He has received the Dick Volz Best 2007 US Ph.D. Thesis Runner-Up Award, RSS - Early Career Spotlight, INNS Young Investigator Award, and IEEE Robotics & Automation Society’s Early Career Award, as well as numerous best paper awards. He received an ERC Starting Grant and was appointed an IEEE fellow, AIAA fellow, and ELLIS fellow.